What is the electric flux through a sphere of radius 4m that contains a (a)



≠+50 hC and (b) -50 hC charge at its center?

1
Expert's answer
2022-02-24T09:05:52-0500

The Gauss's law says

"{\rm flux}=\oint EdA=\frac{q}{\epsilon_0}"

(a)

"{\rm flux}=\frac{50*10^{-9}\:\rm C}{8.85*10^{-12}\:\rm F/m}=5.65*10^3\:\rm N\cdot m^2\cdot C^{−1}"

(b)

"{\rm flux}=\frac{-50*10^{-9}\:\rm C}{8.85*10^{-12}\:\rm F/m}=-5.65*10^3\:\rm N\cdot m^2\cdot C^{−1}"